563blackghost (563blackghost):

Given triangle ABC with A(-4,-2), and B(4,4), and C (18,-8). Write the equation of the line containing the altitude that passes through B in standard form.

563blackghost (563blackghost):

@jim_thompson5910 is there any chance you know how to do this question and can help me with it like finding the altitude?

jimthompson5910 (jim_thompson5910):

You'll first need to find the slope of the line through point A and point C

jimthompson5910 (jim_thompson5910):

slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

tell me what slope you get

563blackghost (563blackghost):

I got 3/11

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):

now to find the perpendicular slope, we do 2 things a) flip the fraction: 3/11 ----> 11/3 b) flip the sign (positive to negative): 11/3 ----> -11/3

jimthompson5910 (jim_thompson5910):

the original slope is 3/11 the perpendicular slope is -11/3

jimthompson5910 (jim_thompson5910):

we know this altitude has a slope of -11/3 and it goes through B(4,4)

jimthompson5910 (jim_thompson5910):

y = mx+b y = (-11/3)*x + b ... plug in m = -11/3 4 = (-11/3)*(4) + b ... plug in (x,y) = (4,4) now solve for b

563blackghost (563blackghost):

Thanx so much I got it from here xD the confusing part for me was getting the slope and the points but I now understand...thanx again ^^

jimthompson5910 (jim_thompson5910):

you're welcome